AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME  ELOG logo
Message ID: 455     Entry time: Tue Oct 17 08:43:47 2023
Author: CB 
Subject: NEG activation -2 

09:40 - Found temperature and pressure stablised, but too high to comfortably activate NEGs closest to detectors. Detector temperature around 90C. This was expected. Ramped down UHV1400 #1,2,3,4 to P=50W (200C?). Ramped down heaters to 100, 80, 100C. Ramped down heating jacket to 60C.

10:00 - Activated set #5 (x2) at 80W tramp=30, thold=infinity; set#6 (x3) at 120W rest as #5. This is different from previous procedure because we have 4x DSSD but sensor only one one of them. It seems less dangerous to heat up everything at once so that the one TC we have is hopefully representative of the situation.

10:15 - CARME cooling. No visible effects on temperature. Tried P#5 100, P#6 150 W. (500C roughly?) Will probably have to decrease.

10:25 - Detectors getting hotter quickly, but still below 100C. No changes for now.

10:35 - As expected, detectors are getting too hot, too quickly. Decreased power to 60, 90 W

10:43 - Ramped down #5,6 to 40, 60 W. Should be around 200C at equilibrium.

10:45 - Also decreased #1,2,3,4 to 40W, and D2000 to 150C.

10:52 - Further decreased #5,6 to 10, 15W.

11:04 - Temperature back under control

11:10 - Detectors actually cooling now. Increased #5,6 to 20,30 W.

11:18 - Detectors still cooling. Increased #5.6 to 30,45W. If stable plan to leave like this until later this afternoon.

13:23 - Temperatures stable. Increased #5,6 to 40, 60 W. Decreased heaters to 80,80,80 C. Decreased jacket to 40C. Opened the tent partly to connect the vacuum division non-bakeable ion pump cable.

13:35 - Ion pump cable connected and tent closed as best I can. Interestingly, the pressure is going up probably as a result of the detector/kapton cables temperature going up due to the slight increase in #5,6 power.

14:15? - Graziano S. flashed the vacuum divison ion pump.

14:25 - Pressure keeps increasing slightly. I am taking this as a sign the kapton is outgassing, which is being made to out gass which is good. Decided to leave temperatures as they are until later today and start cooling down in the late afternoon / evening. Decreased slightly #5,6 to 34, 51W. Increased D2000 to 200C. Will leave like this if stable.

15:00 - Looks stable. See attachment 3. Leaving for a few hours, before proceeding to gradual cool down overnight. Will ask Jan for help in turning everything really off.

18:00 - All NEGs off. All apadev 10&11 channels to 40C. Ion pumps remain off. Asked JG to come tomorrow late morning and turn the apadev channels off and the ions pumps on.

Attachment 1: Screenshot_2023-10-17_09-45-23.png  219 kB  Uploaded Tue Oct 17 09:45:31 2023  | Hide | Hide all
Screenshot_2023-10-17_09-45-23.png
Attachment 2: Screenshot_2023-10-17_11-04-09.png  206 kB  Uploaded Tue Oct 17 11:04:19 2023  | Hide | Hide all
Screenshot_2023-10-17_11-04-09.png
Attachment 3: Screenshot_2023-10-17_14-56-55.png  205 kB  Uploaded Tue Oct 17 14:57:20 2023  | Hide | Hide all
Screenshot_2023-10-17_14-56-55.png
Attachment 4: Screenshot_2023-10-17_17-58-24.png  208 kB  Uploaded Tue Oct 17 17:58:32 2023  | Hide | Hide all
Screenshot_2023-10-17_17-58-24.png
Attachment 5: Screenshot_2023-10-17_17-58-51.png  347 kB  Uploaded Tue Oct 17 17:59:05 2023  | Hide | Hide all
Screenshot_2023-10-17_17-58-51.png
ELOG V3.1.4-unknown